How many moles are in 49.7 grams of BaO?


1
Expert's answer
2021-03-05T03:31:45-0500

Solution:

mol(BaO)=137+16=153g/mol

Xmol=49,7/153=0.32483

Xmol=0.32483mol

Answer:mol(BaO)=0.32483mol